# Efficient designs for multivariate crossover trials

### v1

@@ Abstract
This article aims to study efficient/trace optimal designs for crossover trials, with multiple response recorded from each subject in each time period. A multivariate fixed effect model is proposed with direct and carryover effects corresponding to the multiple responses and the error dispersion matrix allowing for correlations to exist between and within responses. Two correlation structures, namely the proportional and the generalized markov covariances are studied. The corresponding information matrices for direct effects under the two covariances are used to determine efficient designs. Efficiency of orthogonal array designs of Type $I$ and strength $2$ is investigated for the two covariance forms. To motivate the multivariate crossover designs, a gene expression data in a $3\times 3$ framework is utilized.
@@ 1 Introduction
In many biomedical and clinical studies, we frequently come across crossover or repeated measure setups with $g$ responses, where $g>1$. For instance, Grender and Johnson, (1993) studies the effect of caffeine on the mathematical ability of boys aged 14 and 16 years in a crossover framework. In this study, two responses diastolic and systolic blood pressures, of each boy are recorded in each of the two periods in a $2\times 2$ crossover design. More recently, Leaker et al., (2017) considered a randomized, placebo controlled, double-blind $3\times 3$ crossover trial, where in each period gene expression values corresponding to multiple genes were recorded for each subject, resulting in a multiple response crossover trial. Due to more than one response being measured from the same subject at same and different time periods, there is a necessity to account for both between and within response correlations in the covariance structure. In this article, our objective is to find trace optimal designs in a $p$-period and $t$- treatment crossover setup for $g$ correlated responses. Though these correlated response crossover trials are popular in the statistical literature (see Grender and Johnson, 1993; Chinchilli and Esinhart, 1996; Tudor et al., 2000; Pareek et al., 2023), mostly estimation and testing of hypothesis for such trials have been studied without any attention being paid to the actual designing of such studies. To the best of our knowledge this is the first work in optimal designs for such correlated response crossover experiments.
@@
The key difficulty in finding optimal designs for such crossover models is specifying the direct and cross covariance for the multiple responses. In the direct covariance, correlations arising from within responses measured in different time periods from each subject are acknowledged. This closely resembles the covariance pattern we see in usual longitudinal/repeated measures studies. To consider the relations between different responses measured in the same and different time periods we use cross correlations. Thus resembling a cokriging spatial experiment, only instead of spatial units, here we have time periods. In this article, we propose a multivariate response model for a $p,t\geq 3$ crossover setup with correlation within and between responses. Two popular covariance structures from the statistical literature, mainly proportional covariance and the markovian-type covariance (see Journel, 1999; Chiles and Delfiner, 2012; Dasgupta et al., 2022), are utilized to model the correlations within (direct covariance) and between (cross covariance) responses. As in all crossover studies, subjects are assumed to be uncorrelated. Information matrices for the proportional and markovian structures are determined and later used to obtain optimal designs. We show that these information matrices fail to satisfy the usual sufficient conditions (Kiefer, 1975) and the conditions by Yeh, (1986) for universal optimal designs, leading us to instead search for trace optimal or efficient crossover designs (Bate and Jones, 2006). For the proportional structure, an orthogonal array design represented by $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$, is proved to be trace optimal/efficient. Under the more complicated markovian structure, for $g=2$, i.e., a bivariate model, we explore high efficiency design for various choices of within and between response covariance matrices.
@@
Various researchers namely, Matthews, (1987, 1990), Kunert, (1991), Stufken, (1991), Kushner, 1997b ; Kushner, (1998), Bate and Jones, (2006) and Singh and Kunert, (2021) have studied efficient/highly efficient crossover designs. While, universally optimal crossover designs have been discussed by Hedayat and Afsarinejad, (1978), Cheng and Wu, (1980), Kunert, (1984), Kunert and Martin, (2000), Hedayat and Yang, (2003, 2004), to name a few. For a detailed review of crossover designs, we would like to refer to the paper by Bose and Dey, (2013) and books by Senn, (2002), Bose and Dey, (2009) and Kenward and Jones, (2014). However, all of these works concentrate only on single responses measured in each period. Very recently, Niphadkar and Mukhopadhyay, (2023) have determined universal optimality results for multiple response crossover trials, where the responses are assumed to be uncorrelated with a homoscedastic covariance matrix.
@@
The original contributions of this article include (i) a multivariate response crossover model with proportional and markovian covariance structures to model correlations, (ii) information matrices for direct effects, (iii) trace optimal design under the proportional covariance structure and (iv) highly efficient designs for the markovian structure.
@@ 1.1 Motivating example: A gene expression study
We motivate our multivariate crossover model using a gene expression dataset from Leaker et al., (2017). In this study, expressions from multiple genes from the subjects are measured in a randomized, double-blind $3\times 3$ crossover setup. The main aim of the study was to understand the biomarkers of mucosal inflammation and transcriptomics. Two doses of an oral drug were compared with a placebo, thus resulting in a $3$ treatment trial. Each of the subjects were assigned to any one of the $3$ treatment sequences; $ABC$, $CAB$ and $BCA$, where treatment $A$ represented the $10$ mg dose of the drug, while $B$ and $C$ were the placebo and $25$ mg dose of the drug, respectively. In this study cytokines, chemokines, and nasal mucosal gene expressions were analyzed with main focus on type $2$ inflammation related to IL-$5$ and IL-$13$. For the purpose of illustration, we considered measurements from $10$ genes in each of the periods. The information about these $10$ genes is given in Table S.1. The dataset on gene profile is publicly available at the NCBI Gene Expression Omnibus (Clough, 2016) with accession number $GSE67200$. Figures 0(a) and 0(c) display the plot of average response values for each gene, where the averaging is done over subjects. Figure 0(b) displays similar plot with averaging done over periods. From Figures 0(a), 0(b) and 0(c) we note that effects due to periods, subjects and treatments differ across genes. In Figure 0(d) we plot the average response values corresponding to each gene. From the varying height of the line segments we conclude that the mean effect for each gene is distinct. We also studied the $10$ gene profiles in details and noted significant correlations between multiple gene pairs.
Thus in the following section we propose a multivariate model that has varying effects corresponding to each response variable and include correlation between and within responses.
@@
Figure 1: Plotting the period, subject, treatment and overall mean corresponding to the $10$ genes.
@@ 2 A multivariate crossover model
We consider the class, $\varOmega_{t,n,p}$, of crossover designs with $t$ treatments, $n$ subjects and $p$ periods, where $t,p>2$. It is assumed that measurements are recorded on $g\geq 1$ response variables from every subject in each period. Suppose for a design $d\in\varOmega_{t,n,p}$, $Y_{dijk}$ represents the $k^{th}$ response measurement from the $j^{th}$ subject in the $i^{th}$ period of the experiment, where $1\leq k\leq g$, $1\leq i\leq p$ and $1\leq j\leq n$. We model $Y_{dijk}$ as:
@@
where corresponding to the $k^{th}$ response variable, $\mu_{k}$ is the intercept effect, $\alpha_{i,k}$ is the $i^{th}$ period effect, $\beta_{j,k}$ is the $j^{th}$ subject effect, $\tau_{s,k}$ and $\rho_{s,k}$ are respectively the direct effect and the first order carryover effect due to the $s^{th}$ treatment where $1\leq s\leq t$, and $d(i,j)$ denotes the treatment allocated to the $j^{th}$ subject in the $i^{th}$ period. We assume that (2.1) is a fixed effect model with a zero mean error term, $\epsilon_{ijk}$, and no carryover effect in the first period. If $g=1$, then the above model becomes an univariate response model similar to those found in the statistical literature (see Bose and Dey, 2009).
@@
$\begin{bmatrix}\bm{Y}^{{}^{\prime}}_{d1}&\ldots&\bm{Y}^{{}^{\prime}}_{dg}\end{bmatrix}^{{}^{\prime}}$ $=\bm{Z}_{d}\bm{\eta}+\begin{bmatrix}\bm{\epsilon}^{{}^{\prime}}_{1}&\ldots&\bm{\epsilon}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}},$ (2.4)
where $\bm{Y}_{dk}=\left(Y_{d11k},\ldots,Y_{dp1k},\ldots,Y_{d1nk},\ldots,Y_{dpnk}\right)^{{}^{\prime}}$ and $\bm{\epsilon}_{k}=\left(\epsilon_{11k},\ldots,\epsilon_{pnk}\right)^{{}^{\prime}},$ represent the observations and the error terms, respectively, corresponding to the $k^{th}$ response, and $\bm{\eta}=$
$\begin{bmatrix}\mu_{1}&\ldots&\mu_{g}&\bm{\alpha}^{{}^{\prime}}_{1}&\bm{\beta}^{{}^{\prime}}_{1}&\ldots&\bm{\alpha}^{{}^{\prime}}_{g}&\bm{\beta}^{{}^{\prime}}_{g}&\bm{\tau}^{{}^{\prime}}_{1}&\bm{\rho}^{{}^{\prime}}_{1}&\ldots&\bm{\tau}^{{}^{\prime}}_{g}&\bm{\rho}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}}$ is the parameter vector. The design matrix is given by,
@@
where $\bm{Z}_{1}=\bm{I}_{g}\otimes\bm{X}_{1}$, $\bm{Z}_{2}=\bm{I}_{g}\otimes\bm{X}_{2}$, $\bm{X}_{1}=\begin{bmatrix}\bm{P}&\bm{U}\end{bmatrix}$ and $\bm{X}_{2}=\begin{bmatrix}\bm{T}_{d}&\bm{F}_{d}\end{bmatrix}$. For the periods and subjects, we have $\bm{P}=\bm{{1}}_{n}\otimes\bm{I}_{p}$, $\bm{U}=\bm{I}_{n}\otimes\bm{{1}}_{p}$, while $\bm{T}_{d}$ and $\bm{F}_{d}$ are the matrices for treatment and carryover effects, respectively. Note that $\bm{F}_{d}=\left(\bm{I}_{n}\otimes\bm{\psi}\right)\bm{T}_{d}$, where $\bm{\psi}=\begin{bmatrix}\bm{{0}}^{{}^{\prime}}_{p-1\times 1}&0\\ \bm{I}_{p-1}&\bm{{0}}_{p-1\times 1}\end{bmatrix}$. The vector of errors is assumed to have $E(\bm{\epsilon})=\bm{{0}}_{gnp}$, with dispersion
$\mathbb{D}\left(\bm{\epsilon}\right)$ $=\bm{\varSigma}=\begin{bmatrix}\bm{\varSigma}_{11}&\ldots&\bm{\varSigma}_{1g}\\ \vdots&\vdots&\vdots\\ \bm{\varSigma}_{g1}&\ldots&\bm{\varSigma}_{gg}\end{bmatrix},$ (2.10)
where $\bm{\varSigma}_{k^{\prime}k}=\bm{\varSigma}_{kk^{\prime}}^{{}^{\prime}}$, for $k\neq k^{\prime}=1,\ldots,g$. Here, the symmetric matrix $\bm{\varSigma}_{kk}$ represent the variance of $\bm{\epsilon}_{k}$, while $\bm{\varSigma}_{kk^{\prime}}$ represents the covariance between $\bm{\epsilon}_{k}$ and $\bm{\epsilon}_{k^{\prime}}$. The matrix $\bm{\varSigma}$ is assumed to be positive definite. Also, observations from different subjects are taken to be uncorrelated, as is the usual practice (see Bose and Dey, 2009).
Two different structures are considered to model the covariances between responses, namely, the proportional structure and the generalized markov-type structure. The generalized markov-type structure for the spatial data was proposed recently by Dasgupta et al., (2022) and is a generalization of the markov-type structure discussed in Journel, (1999) and Chiles and Delfiner, (2012). The proportional structure is simpler than the markovian one and is used for the general $g>1$ case. However, in the statistical literature, the markovian structure has been studied for only the $g=2$ case. We also restrict our search for optimal designs for a markovian structure to $g=2$.
@@
Under the proportional structure the dispersion matrix $\bm{\varSigma}$ is as follows:
$\bm{\varSigma}$ $=\bm{\varGamma}\otimes\left(\bm{I}_{n}\otimes\bm{V}\right),$
where $\bm{\varGamma}$ is assumed to be a $g\times g$ positive definite and symmetric matrix having non-zero off diagonal elements, and $\bm{V}$ a known positive definite and symmetric matrix. Under the proportional structure, the within and between responses have similar covariance matrices differing by some constants.

### v2

@@ Abstract
This article aims to study efficient/trace optimal designs for crossover trials with multiple responses recorded from each subject in the time periods. A multivariate fixed effects model is proposed with direct and carryover effects corresponding to the multiple responses. The corresponding error dispersion matrix is chosen to be either of the proportional or the generalized Markov covariance type, permitting the existence of direct and cross-correlations within and between the multiple responses. The corresponding information matrices for direct effects under the two types of dispersions are used to determine efficient designs. The efficiency of orthogonal array designs of Type $I$ and strength $2$ is investigated for a wide choice of covariance functions, namely, Mat($0.5$), Mat($1.5$) and Mat($\infty$). To motivate these multivariate crossover designs, a gene expression dataset in a $3\times 3$ framework is utilized.
@@ 1 Introduction
In biomedical and clinical studies, we often come across crossover or repeated measure setups with $g$ responses, where $g>1$. For instance, Grender and Johnson, (1993) studies the effect of caffeine on the mathematical ability of boys aged 14 and 16 years in a crossover framework. In this study, two responses diastolic and systolic blood pressures, of each boy are recorded from each of the two periods in a $2\times 2$ crossover design setup. More recently, Leaker et al., (2017) considered a randomized, placebo controlled, double-blind $3\times 3$ crossover trial, where in each time period gene expression values corresponding to multiple genes were recorded for each subject, resulting in a multiple response crossover trial. Due to more than one response being measured from the same subject at the same and different time periods, there is a necessity to account for both between and within response correlations in the covariance structure. Using the gene experiment study of Leaker et al., (2017), we motivate a multivariate response model for a $p,t\geq 3$ crossover setup with correlation within and between responses. Our main objective is to find $p$-period and $t$-treatment trace optimal/efficient crossover designs for $g$ correlated responses. Though these correlated response crossover trials are quite popular in the statistical literature (see Grender and Johnson, 1993; Chinchilli and Esinhart, 1996; Tudor et al., 2000; Pareek et al., 2023), mostly the topics of estimation and testing of hypothesis for such trials have been studied without any attention being paid to the actual designing of such studies. To the best of our knowledge, this is the first work in optimal designs for correlated response crossover experiments.
@@
The key difficulty in finding optimal designs for such crossover models is specifying the direct and cross-covariance for the multiple responses. In the direct covariance, correlations arising from within responses measured in different time periods from each subject are acknowledged. This closely resembles the covariance pattern we see in usual longitudinal/repeated measures studies. To consider the relations between different responses measured in the same and different time periods we use cross-correlations. Two popular covariance structures from the statistical literature, resembling the proportional covariance and the generalized Markovian-type covariance (see Journel, 1999; Chiles and Delfiner, 2012; Dasgupta et al., 2022), are utilized to model the correlations within (direct covariance) and between (cross-covariance) responses. Under the proportional structure, the between response covariance is assumed to be proportional to the within response covariance, $\bm{V}$, where $\bm{V}$ is positive definite. The Markov-type covariance has been proposed in the statistical literature solely for $g=2$ or the bivariate response case, with the within and between response covariances formed by two respective matrices, $\bm{V}_{1}$ and $\bm{V}_{R}$, where $\bm{V}_{1}$ is a positive definite matrix and $\bm{V}_{R}$ is a valid correlation matrix. Possible choices of covariance functions, for example the Exponential (Mat($0.5$)), Gaussian (Mat($\infty$)) and Matern are used in this article for $\bm{V}$ and ($\bm{V}_{1}$, $\bm{V}_{R}$) (see Section 2.1 for details). Thus, it is important to note that the proportional and Markovian-type covariances are not at all restrictive in nature, in fact they allow us to explore a wide range of covariance functions for modelling the within and between response correlations.
For obtaining optimal designs, information matrices under the proportional and generalized Markovian-type covariance assumptions are determined. We show that these information matrices fail to satisfy the usual sufficient conditions (Kiefer, 1975) and the conditions by Yeh, (1986) for universal optimal designs, leading us to instead search for trace optimal or efficient crossover designs (Bate and Jones, 2006). For the proportional structure, an orthogonal array design represented by $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$, is proved to be trace optimal/efficient. Under the more complicated Markovian structure, for $g=2$, i.e., a bivariate model, we show that the same orthogonal array design is highly efficient for various choices of within and between response covariance matrices. Further, we utilize the gene experiment dataset to illustrate the choice of a highly efficient design in a multivariate crossover setup. To reiterate, it is very important to understand that in this article we are able to determine efficient designs theoretically, for a wide range of covariance structures as opposed to obtaining optimal designs only for the exponential covariance case (Dasgupta et al., 2022).
@@
Various researchers, namely, Matthews, (1987, 1990), Kunert, (1991), Stufken, (1991), Kushner, 1997b ; Kushner, (1998), Bate and Jones, (2006) and Singh and Kunert, (2021) have studied efficient/highly efficient crossover designs. While, universally optimal crossover designs have been discussed by Hedayat and Afsarinejad, (1978), Cheng and Wu, (1980), Kunert, (1984), Kunert and Martin, (2000), Hedayat and Yang, (2003, 2004), to name a few. For a detailed review of crossover designs, we would like to refer to the paper by Bose and Dey, (2013) and books by Senn, (2002), Bose and Dey, (2009) and Kenward and Jones, (2014). However, all of these works concentrate only on single responses measured in each period. Very recently, Niphadkar and Mukhopadhyay, (2023) have determined universal optimality results for multiple response crossover trials, where the responses are assumed to be uncorrelated with a homoscedastic covariance matrix.
@@
The original contributions of this article include (i) a multivariate response crossover model with proportional and Markovian covariance structures to model correlations, (ii) information matrices for direct effects, (iii) theoretical results for efficient designs for a wide range of covariance functions, particularly -trace optimal design under the proportional covariance structure and highly efficient designs for the Markovian structure.
@@ 1.1 Motivating example: A gene expression study
We motivate our multivariate crossover model using a gene expression dataset from Leaker et al., (2017). In this study, multiple gene expressions from subjects were measured in a randomized, double-blind $3\times 3$ crossover setup to understand the biomarkers of mucosal inflammation. Two doses of an oral drug were compared with a placebo, thus resulting in a $3$ treatment trial. Each of the subjects were assigned to any one of the $3$ treatment sequences; $ABC$, $CAB$ and $BCA$, where treatment $A$ represented the $10$ mg dose of the drug, while $B$ and $C$ were the placebo and $25$ mg dose of the drug, respectively. For the purpose of illustration, we considered responses/measurements corresponding to $3$ genes from each of the $3$ periods, thus in the gene example, $g=3$. The information about these $3$ genes is given in Table S.1. This dataset on gene profile is publicly available at the NCBI Gene Expression Omnibus (Clough and Barrett, 2016) with accession number $GSE67200$. 1(a) and 1(c) display the plot of average response values for each gene with respect to periods and treatments, respectively, where the averaging is done over subjects. 1(b) displays a similar average response plot with respect to subject, where averaging is done over periods. From 1(a), 1(b) and 1(c) we note that effects due to periods, subjects and treatments differ across genes. 1(d) shows that the average response values vary across genes. These plots thus help us to develop a multivariate crossover model in Section 2 with varying intercepts for each gene and differential effects of periods, treatments and subjects on the $3$ gene responses.
We also studied the correlation profile of the $3$ genes in detail and noted significant correlations between multiple gene pairs. In order to check for correlation within and between the $3$ responses, several tests were performed. The test results are shown in Fig. 2 and Fig. 3. We note from Fig. 2 that at the $5\%$ level of significance, within gene correlation for genes $1$ and $2$ are high between periods, while in Fig. 3 we see significant correlations present between gene pairs $(1,2)$ and $(2,3)$.
@@
Figure 1: Plots of the period, subject, treatment and gene versus average response corresponding to the $3$ genes.
(a) $Gene_{1}$
(b) $Gene_{2}$
(c) $Gene_{3}$
Figure 2: Within response correlation plots corresponding to the $3$ genes. The values are the p-values of the corresponding correlation tests while the colours represent the quantity of correlation present.
(a) $\left(Gene_{1},Gene_{2}\right)$
(b) $\left(Gene_{1},Gene_{3}\right)$
(c) $\left(Gene_{2},Gene_{3}\right)$
Figure 3: Between response correlation plots corresponding to the $3$ gene pairs. The values are the p-values of the corresponding correlation tests while the colours represent the quantity of correlation present.
Learning from the motivating gene example, in the following section we propose a multivariate model that has varying effects of periods, subjects and treatments corresponding to each response variable, and also includes correlation between and within responses.
@@ 2 A multivariate crossover model
We consider the class, $\Omega_{t,n,p}$, of crossover designs with $t$ treatments, $n$ subjects and $p$ periods, where $t,p\geq 3$. It is assumed that measurements are recorded on $g\geq 1$ response variables from every subject in each period. Suppose for a design $d\in\Omega_{t,n,p}$, $Y_{dijk}$ represents the $k^{th}$ response measurement from the $j^{th}$ subject in the $i^{th}$ period of the experiment, where $1\leq k\leq g$, $1\leq i\leq p$ and $1\leq j\leq n$. We model $Y_{dijk}$ as:
@@
where corresponding to the $k^{th}$ response variable, $\mu_{k}$ is the intercept effect, $\alpha_{i,k}$ is the $i^{th}$ period effect, $\beta_{j,k}$ is the $j^{th}$ subject effect, $\tau_{s,k}$ and $\rho_{s,k}$ are respectively the direct effect and the first order carryover effect due to the $s^{th}$ treatment where $1\leq s\leq t$, and $d(i,j)$ denotes the treatment allocated to the $j^{th}$ subject in the $i^{th}$ period. The assumptions of varying effects are in accordance with the observations made in the gene expression study in Subsection 1.1. We assume that (2.1) is a fixed effect model with a zero mean error term, $\epsilon_{ijk}$, and no carryover effect in the first period. If $g=1$, then the above model becomes an univariate response model similar to those found in the statistical literature (see Bose and Dey, 2009).
@@
$\begin{bmatrix}\bm{Y}^{{}^{\prime}}_{d1}&\cdots&\bm{Y}^{{}^{\prime}}_{dg}\end{bmatrix}^{{}^{\prime}}$ $=\bm{Z}_{d}\bm{\eta}+\begin{bmatrix}\bm{\epsilon}^{{}^{\prime}}_{1}&\cdots&\bm{\epsilon}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}},$ (2.2)
where $\bm{Y}_{dk}=\left(Y_{d11k},\cdots,Y_{dp1k},\cdots,Y_{d1nk},\cdots,Y_{dpnk}\right)^{{}^{\prime}}$ and $\bm{\epsilon}_{k}=\left(\epsilon_{11k},\cdots,\epsilon_{pnk}\right)^{{}^{\prime}},$ represent the observations and the error terms, respectively, corresponding to the $k^{th}$ response, and $\bm{\eta}=$
$\begin{bmatrix}\mu_{1}&\cdots&\mu_{g}&\bm{\alpha}^{{}^{\prime}}_{1}&\bm{\beta}^{{}^{\prime}}_{1}&\cdots&\bm{\alpha}^{{}^{\prime}}_{g}&\bm{\beta}^{{}^{\prime}}_{g}&\bm{\tau}^{{}^{\prime}}_{1}&\bm{\rho}^{{}^{\prime}}_{1}&\cdots&\bm{\tau}^{{}^{\prime}}_{g}&\bm{\rho}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}}$ is the parameter vector. The design matrix is given by
@@
where $\bm{Z_{1}}=\bm{I}_{g}\otimes\bm{X_{1}}$, $\bm{Z_{2}}=\bm{I}_{g}\otimes\bm{X_{2}}$, $\bm{X_{1}}=\begin{bmatrix}\bm{P}&\bm{U}\end{bmatrix}$ and $\bm{X_{2}}=\begin{bmatrix}\bm{T}_{d}&\bm{F}_{d}\end{bmatrix}$. For the periods and subjects, we have $\bm{P}=\bm{{1}}_{n}\otimes\bm{I}_{p}$, $\bm{U}=\bm{I}_{n}\otimes\bm{{1}}_{p}$, while $\bm{T}_{d}$ and $\bm{F}_{d}$ are the matrices for treatment and carryover effects, respectively. Note that $\bm{F}_{d}=\left(\bm{I}_{n}\otimes\bm{\psi}\right)\bm{T}_{d}$, where $\bm{\psi}=\begin{bmatrix}\bm{{0}}^{{}^{\prime}}_{p-1\times 1}&0\\ \bm{I}_{p-1}&\bm{{0}}_{p-1\times 1}\end{bmatrix}$. The vector of errors is assumed to have multivariate normal distribution with $E(\bm{\epsilon})=\bm{{0}}_{np\times 1}$ and the dispersion matrix as
$\mathbb{D}\left(\bm{\epsilon}\right)$ $=\bm{\mathit{\Sigma}}=\begin{bmatrix}\bm{\mathit{\Sigma}}_{11}&\cdots&\bm{\mathit{\Sigma}}_{1g}\\ \vdots&\vdots&\vdots\\ \bm{\mathit{\Sigma}}_{g1}&\cdots&\bm{\mathit{\Sigma}}_{gg}\end{bmatrix},$ (2.4)
where $\bm{\mathit{\Sigma}}_{k^{\prime}k}=\bm{\mathit{\Sigma}}_{kk^{\prime}}^{{}^{\prime}}$, for $k\neq k^{\prime}=1,\cdots,g$. Here, the symmetric matrices, $\bm{\mathit{\Sigma}}_{kk}$ represent the variance of $\bm{\epsilon}_{k}$, while $\bm{\mathit{\Sigma}}_{kk^{\prime}}$ represents the covariance between $\bm{\epsilon}_{k}$ and $\bm{\epsilon}_{k^{\prime}}$. The matrix $\bm{\mathit{\Sigma}}$ is assumed to be positive definite. Also, observations from different subjects are taken to be uncorrelated, as is the usual practice (see Bose and Dey, 2009).
The investigation made regarding the presence of correlation between genes (see Fig. 3) in Subsection 1.1, motivates the use of the covariance structure with cross-covariances. Two different structures are considered to model the covariances between responses, namely, the proportional structure and the generalized Markov-type structure. A generalized version of the Markov-type covariance structure for spatial data (Journel, 1999; Chiles and Delfiner, 2012) was proposed recently by Dasgupta et al., (2022). The proportional structure is simpler than the Markovian one and is used for the general $g>1$ case. However, in the statistical literature, the Markovian structure has been studied for only the $g=2$ or bivariate response case. We also restrict our search for optimal designs for a Markovian covariance structure to $g=2$ or bivariate responses.
@@
Under the proportional structure the dispersion matrix $\bm{\mathit{\Sigma}}$ is taken as follows:
$\bm{\mathit{\Sigma}}$ $=\bm{\mathit{\Gamma}}\otimes\left(\bm{I}_{n}\otimes\bm{V}\right),$
where $\bm{\mathit{\Gamma}}$ is assumed to be a $g\times g$ positive definite and symmetric matrix having non-zero off-diagonal elements, and $\bm{V}$ a known positive definite and symmetric matrix. For $g=2$ case, the matrix $\bm{\mathit{\Gamma}}$ is given as
$\bm{\mathit{\Gamma}}=\begin{bmatrix}\gamma_{11}&\gamma_{12}\\ \gamma_{12}&\gamma_{22}\end{bmatrix},$
where $\gamma_{11},\gamma_{22}>0$ and $\gamma_{11}\gamma_{22}>\gamma_{12}^{2}$. Under the proportional structure, the within and between responses have similar covariance matrices differing by some constants.